# Deep Unlearn: Benchmarking Machine Unlearning

### v1

@@ 1 Introduction
Machine unlearning aims to remove the influence of a specified subset of training data points from trained models [10]. This process is crucial for enhancing privacy preservation, model safety, and overall model quality. MU helps ensure compliance with the right to be forgotten [23], removes erroneous data points that negatively impact model performance [43], and eliminates biases introduced by parts of the training data [14]. Deep neural networks present significant challenges for MU due to their computationally intensive training requirements, highly non-convex loss landscapes, and tendency to memorize substantial portions of their training data [22, 12, 21]. The key open challenges in MU include: (1) A degradation in model accuracy often accompanies unlearning; (2) Some MU methods require the model to be trained in specific ways, such as saving checkpoints, tracking accumulated gradients, or training with differential privacy, limiting their applicability to already deployed models; (3) Assurance of information removal is difficult as there are no reliable metrics to measure it accurately; (4) There is no consensus on which methods are the most effective.
@@
Q1. Are the commonly used MU baselines reliable? Most recently introduced MU methods are compared only with three baselines: Fine-tuning, Gradient Ascent, and Successive Random Labels, and not across recently introduced approaches. We show that with proper hyperparameter selection FT is a reliable baseline. In contrast, GA consistently performs poorly and should be replaced by the more recent Negative Gradient Plus (NG+) [37] that simultaneously reduces the performance in the forget set while maintaining the performance on the rest of the data points.
Q2. How reliable are MU methods across datasets, models, and initializations? Our findings show that, unlike most recent MU methods, Masked-Small-Gradients is consistently among the best performing methods across various metrics. In contrast, methods such as SRL do not consistently outperform others across different datasets.
@@ 2 Background of Machine Unlearning
Setting. Starting with a training set $\mathcal{D}=\{(\mathbf{x}_{i},y_{i})\}_{i=1}^{N}$ and a trained model $f_{O}$, referred to as the original model, the objective of a MU method $U$ is to remove the influence of a particular subset of training set $\mathcal{D}_{F}=\{(\mathbf{x}_{i},y_{i})\}_{i=1}^{K}\subset\mathcal{D}$ referred to as the forget set where $K\ll N$. The rest of training set $\mathcal{D}_{R}=\mathcal{D}\backslash\mathcal{D}_{F}$ is called the retain set. The forget set and the retain set are distinct and complementary subsets of the training set. The outcome of MU is an unlearned model $f_{U}$, the aim for which is to perform on par with a model retrained from scratch on $\mathcal{D}_{R}$; this latter model $f_{R}$ is referred to as the retrained model. We denote the weights of the original model and the retrained model as $\theta_{O}$ and $\theta_{R}$, respectively. For evaluations, we consider two held-out sets: the validation set $\mathcal{D}_{V}$ and test set $\mathcal{D}_{T}$, both drawn from the same distribution as $\mathcal{D}$. We consider the accuracy of the retrained model as the optimal accuracy. One critical assumption we make is that the MU method has access to $\theta_{O}$, $\mathcal{D}_{R}$, $\mathcal{D}_{F}$, and $\mathcal{D}_{V}$.
@@
(1) Unlearning Evaluation. To evaluate the success of unlearning, one approach is to check whether data points in $\mathcal{D}_{F}$ still influence the predictions made by the unlearned model  [13, 37, 31]. This is commonly done via influence functions [35, 8, 29], membership inference attacks (MIA) [44]. MIA has become one of the most common approaches for evaluating MU methods. It aims to determine whether specific data points were part of the original training dataset based on the unlearned model. We compute the population-based U-MIA, denoted with $\text{MIA}(\mathcal{D},f_{U})$ evaluated on data $\mathcal{D}$ [^1] The methodology for this follows the classic MIA: we compute the losses on $\mathcal{D}_{F}$ and $\mathcal{D}_{V}$, we shuffle and trim them so that they are of equal size. We then train logistic regression models in a 10-fold cross validation, and compute the average accuracy across the folds.. Using this we define the discernibility metric as $\text{Disc}(\mathcal{D}_{V},f_{U})=|2\times\text{MIA}(\mathcal{D},f_{U})-1|\in[0,1]$, and similarly, indiscernibility is given by $\text{Indisc}(\mathcal{D},f_{U})=1-\text{Disc}(\mathcal{D},f_{U})\in[0,1]$. We set $\mathcal{D}$ to either the test set $\mathcal{D}_{T}$ or validation set $\mathcal{D}_{V}$. The indiscernibility equals $1$ when the accuracy of the MIA is not better than random guessing. A more recent MIA variation is the unlearning likelihood ratio attack (U-LiRA) [37]. As highlighted by [31], U-LiRA is a more robust evaluation approach for approximate MU. Nonetheless, U-LiRA is much more computationally demanding than U-MIA. We evaluate the methods that defeat the weaker, less expensive U-MIA attack, additionally against the U-LiRA attack.
@@
(2) Accuracy. The classification accuracy of unlearned model on the retain set should be as close as possible to that of the original model. First, we consider three metrics derived directly from the model’s classification accuracy on different sets: retain accuracy (RA), forget accuracy (FA), and test accuracy (TA). RA is defined as
@@
and formulas for FR and TR can be derived using FA and TA, respectively. An unlearned model with a score of $1$ indicates that its accuracy perfectly matches the accuracy of the reference retrain model. A score below $1$ indicates that the model under-performs and a score above $1$ indicates that the model over-performs. We further define the retention deviation (RetDev) as:
$\text{RetDev}=|RR(f_{U},f_{R})-1|+|FR(f_{U},f_{R})-1|+|TR(f_{U},f_{R})-1|\in[0,+\infty),$ (3)
which provides information on the cumulative divergence of the unlearned model in terms of retention score. The closer to 0, the better as 0 indicates that the model perfectly matches the performance of the retrained model.
(3) Efficiency. run-time efficiency (RTE) of a MU method should ideally be lower than the naive approach of just retraining the model from scratch on the retain set. As the high computational cost of the retraining algorithm motivated the development of approximate MU methods, we evaluate how much faster each MU method is compared to the retraining algorithm. We define RTE of $U$ as the number of seconds it takes to complete, denoted as $\text{RT}(U)$ (we use the same machine and resources for all the experiments). To indicate the relative speedup compared to the retrained model, we define RTE of an unlearn method $U$ as:
@@
where $U_{R}$ denotes the retrain method. The RTE of retraining from scratch would thus be 1; any method with an RTE less than 1 is slower than retraining from scratch, and vice versa.
@@ 3 Machine Unlearning Methods
We briefly discuss the main unlearning methods considered.
@@ 3.1 Classical Baselines
FineTune (FT) finetunes the original model $f_{O}$ on only the retain set $\mathcal{D}_{R}$ for several epochs.
Successive Random Labels (SRL) the model is trained on both the forget set $\mathcal{D}_{F}$ and $\mathcal{D}_{R}$ where the labels of $\mathcal{D}_{F}$ are randomly assigned at each epoch.
Gradient Ascent (GA) trains the model using gradient ascent steps on the $\mathcal{D}_{F}$.
@@ 3.2 State-of-the-art MU methods
Expanding upon the classical baselines, we additionally evaluate 15 recent MU methods. We first discuss the seven top-performing methods from the Machine Unlearning Competition 2023 on Kaggle.
Forget-Contrast-Strengthen (FCS) [1] minimizes the Kullback-Leibler Divergence (KLD) between the model’s output on $\mathcal{D}_{F}$ and a uniform distribution over the output classes, then alternatively optimizes a contrastive loss between the model’s outputs on $\mathcal{D}_{R}$ and $\mathcal{D}_{F}$, and minimizes the cross-entropy loss on $\mathcal{D}_{R}$.

### v2

@@ Abstract
Index Terms:
Machine Unlearning, Deep Learning, Privacy, Machine Learning
@@ 1 Introduction
Figure 1: Pipeline on a single dataset. Phase 1: The dataset is split into the data to retain and forget, and evaluation splits. Phase 2: The Original and Retrained models (Reference) are trained across ten shared model initialization seeds. Phase 3: The Machine Unlearning methods are applied to the Original model after three searches with 100 attempts to determine the best hyper-parameters. Phase 4: The model post-unlearning is evaluated with and without the Retrained model across four aspects: Utility, Privacy, Efficiency, and Reliability.
Machine unlearning aims to remove the influence of a specified subset of training data points from trained models [1]. This process is crucial for enhancing privacy preservation, model safety, and overall model quality. MU helps ensure compliance with the right to be forgotten [2], removes erroneous data points that negatively impact model performance [3], and eliminates biases introduced by parts of the training data [4]. Deep neural networks present significant challenges for MU due to their computationally intensive training requirements, highly non-convex loss landscapes, and tendency to memorize substantial portions of their training data [5, 6, 7].
The key open challenges in MU include: (1) A degradation in model accuracy often accompanies unlearning; (2) Some MU methods require the model to be trained in specific ways, such as saving checkpoints, tracking accumulated gradients, or training with differential privacy, limiting their applicability to already deployed models; (3) Assurance of information removal is difficult as there are no reliable metrics to measure it accurately; (4) There is no consensus on the most effective methods.
@@
  • Q1. Are the commonly used MU baselines reliable? Most recently introduced MU methods are compared only with three baselines: Fine-tuning, Gradient Ascent, and Successive Random Labels, and not across recently introduced approaches. We show that with proper hyperparameter selection, FT is a reliable baseline. In contrast, GA consistently performs poorly and should be replaced by the more recent Negative Gradient Plus (NG+) [14] that simultaneously reduces the performance in the forget set while maintaining the performance on the rest of the data points.
  • Q2. How reliable are MU methods across datasets, models, and initializations? Our findings show that Masked-Small-Gradients is consistently among the best-performing methods across various metrics, unlike most recent MU methods. In contrast, methods such as SRL do not consistently outperform others across different datasets.
@@
We publish the source code used for reproducing the experiments conducted in this paper at https://github.com/xcadet/deepunlearn
2 Related Works
We provide some context on data memorization in Machine Learning (Section 2.1), an overview of Machine Unlearning (MU) (Section 2.2) with a focus on its application to Deep Neural Networks (DNNs) (Section 2.3), the scenario of proceeding to unlearn after the model has been trained (Section 2.4), the link between Machine Unlearning and Differential Privacy (DP) (Section 2.5), and existing benchmarks (Section 2.6).
2.1 Data memorization in Machine Learning
Machine Learning methods aim to learn patterns from their training data. Nonetheless, multiple studies have raised awareness of Machine Learning models’ capacity to leak information about their training datasets.
Ateniese et al. [19] showed that one could use a trained machine learning classifier and obtain information about their training set. They reason that typical ML classifiers, such as Support Vector Machines (SVMs) and Hidden Markov Models (HMMs), must adapt their internal states to extract information from the training data. They showed that trained classifiers could leak information about their training data.
Fredrikson et al. [20] raised awareness of the risk associated with Machine Learning models trained on sensitive data. Their work studied the privacy risks associated with models trained to guide patients’ medical treatment based on sensitive attributes. It showed that the trained model could infer sensitive attributes of patients.
Zhang et al. [21] showed that large neural networks could easily fit randomly labeled data. Since random relabeling breaks the relationship between the image and its actual label, it implies that the effective capacity of large popular neural networks is sufficient for these models to memorize their training set.
Carlini et al. [6] studied the case of generative sequence models that are sometimes trained on sensitive data such as private communication. They discussed the topic of unintended memorization, that is, neural networks can memorize information about the training data unrelated to the task they are trained for. This finding raised significant privacy concerns, especially for models trained on sensitive information, as memorization could occur on data points considered outliers.
2.2 Machine Unlearning
Machine Unlearning is often first associated with the work from Cao et al. [22], followed by Bourtoule et al. [8], which proposes SISA (Sharded, Isolated, Sliced, Aggregated) as an exact unlearning method. For a recent overview of MU, we refer to the survey from Xu et al. [23], which provides a taxonomy of common unlearning methods. Furthermore, Zhang et al. [24] review MU through privacy-preserving and security lenses. The authors cover the Confidentiality, Integrity, and Availability security triad and the need for Data Lineage, which relates to following the movement of data in a machine learning pipeline and understand from where it originates, where it is stored, and how it percolates in the system through transformation. Some might have information on standard MU verification methods, privacy evaluation metrics, and datasets; we defer to the work of Nguyen et al. [25] and Shaik et al. [3].
2.2.1 Data Reorganization
methods focus on directly modifying the data to perform unlearning. It is divided into Data Obfuscation, Data Pruning, and Data Replacement.
  • Data Pruning usually relies on dividing the dataset into multiple sub-datasets and training sub-models on these subsets. This is the category to which SISA [8] relates. Our work does not consider methods associated with this setting as they make assumptions about the training process.
  • Data Replacement attempts to unlearn by replacing the original dataset with transformed data that simplifies unlearning specific samples. For instance, Cao et al. [22] replace the training data with summations of efficiently computable transformations. Like data pruning, these methods tend to make strong assumptions about the training process.
2.2.2 Model Manipulation
methods directly adjust the model parameters to remove the influence of specific data points. Model manipulation is divided into Model Shifting, Model Replacement, and Model Pruning.
  • Model Replacement uses pre-calculated parameters that do not reflect the data to forget to replace parts of the trained model. For instance, when using decision trees, one can replace nodes affected by the forget set by pre-calculated nodes  [29].
2.3 Machine Unlearning for Deep Neural Networks
Initially, MU research primarily focused on linear models such as linear regression and logistic models. Such models allow for the design of methods that assume the convexity of the loss function, rendering them less practical for DNN-based approaches. Since DNNs can memorize parts of their training data, they are particularly relevant targets for MU, even more so when trained on large amounts of potentially personal data. For Deep Learning models, unlearning raises additional challenges: 1) the non-convexity of the loss function of Deep Neural Networks [32], 2) the size of the models inducing large computational costs, 3) the randomness coming from the model’s training process, such as the initialization seed, randomness in the mini-batch generation process, and 4) the fact that any model update impacts subsequent versions of the models, namely the weights at epoch $n+1$ directly depend on the weights at update $n$.
2.4 Post-Hoc Machine Unlearning
Thus, we focus on posthoc MU, a scenario where we assume that the unlearning method is agnostic to the original training process of the model. Under such a scenario, differences exist in terms of data availability at unlearning time. For instance, whether one has access to the original training data $\mathcal{D}$, the retain set $\mathcal{D}_{R}$, the forget set $\mathcal{D}_{F}$, or even some external set such as the validation set $\mathcal{D}_{V}$. Therefore, careful consideration should be given to the data requirement associated with an unlearning method. Indeed, some might require having access to both $\mathcal{D}_{R}$ and $\mathcal{D}_{F}$ at the unlearning time, while others assume that $\mathcal{D}_{R}$ is no longer available [35] making them more practical in real-world scenarios. Throughout our benchmark, we make the same assumption as the NeurIPS2023 Unlearning Challenge [11], where the unlearning methods had access to $f_{O},\mathcal{D}_{R},\mathcal{D}_{F},\mathcal{D}_{V}$
2.5 Machine Unlearning and Differential Privacy
We based our Unlearning definition on Sekhari et al. [36] and refer to their work on the distinction between Differential Privacy and the objective of Machine Unlearning. In a high-level picture, differential privacy is a method for publicly sharing aggregated information about a population by describing the patterns discovered among the groups within the dataset while withholding specific information about individual data points. A randomized algorithm $\mathcal{A}$ is $(\varepsilon,\delta)$-differentially private if for all datasets $D_{1}$ and $D_{2}$ that differ on a single data point, and all $S\subseteq\text{Range}(\mathcal{A})$,
2.6 Machine Unlearning benchmarks
Grimes et al. [37] benchmark methods from the literature using CIFAR-100, a ResNet18 architecture, and an adapted version offline version U-LIRA metric. Their work investigates the effect of iterative unlearning, namely, unlearning occurs in multiple rounds to reflect multiple data unlearning queries.
Li et al. [38] propose a benchmark for unlearning on Large Language Models. They suggest the Weapons of Mass Destruction Proxy (WMDP) dataset composed of $3,668$ multi-choice questions about biosecurity, cybersecurity, and chemical security. They aim to limit autoregressive Language models’ ability to answer queries about hazardous knowledge while maintaining the model’s ability to answer questions about nonhazardous knowledge.
Cheng et al. [39] propose a multi-task multimodal machine unlearning benchmark covering image, text, speech, and video modalities. They follow a retrain-free evaluation framework across nine datasets for different discriminative and generative tasks and modalities. In this work, we focus on image classification and consider the effect of model initializations and the methods at the top of the NeurIPS Machine Unlearning competition.
Ma et al. [40] benchmarks Machine Unlearning with a focus on Copyright Infringement for Text-to-Image Diffusion models. They emphasize the ability of diffusion models to generate outputs that are closely similar to images from their training set. In their work, they consider four viewpoints to address copyright infringement: style—the distinctive patterns of an artist—portrait—an individual’s control over the use of their own portrait—artistic creation figures—for instance, animation characters are often protected by law—and licensed illustrations.
Zhang et al. [41] benchmarks unlearning in Diffusion models. They proposed a dataset consisting of Seed images and stylized versions of these images and focused on evaluating Style and Object unlearning in Diffusion Models.
To our knowledge, Triantafillou et al. [42]’s work is the most similar to ours. Their work leverages the results from the NeurIPS Unlearning Competition and evaluates both the methods from the competition and state-of-the-art methods. We refer to them for additional discussion on the prize-winning methods of the competition. Their work evaluates the methods on CASIA-SURF and FEMNIST using a ResNet-18 and searches for the hyper-parameters using a grid search on CASIA-SURF with three possible values per hyper-parameters We consider five datasets, both Convolutional Neural Networks (CNN) and Vision Transformers (ViT), and perform the hyper-parameter search using the hyper-parameters optimizer framework Optuna [43]. Our findings indicate that the best-performing method from the competition is less reliable when tested across different architectures and datasets. Given that unlearning methods can be sensitive to hyper-parameter choices, our benchmark ensures fair comparisons by searching for 300 hyper-parameter configurations (three searches of 100 combinations each), with each search starting from a model with different initial weights.
@@ 3.1 Problem setting
Starting with a training set $\mathcal{D}=\{(\mathbf{x}_{i},y_{i})\}_{i=1}^{N}$ and a trained model $f_{O}$, referred to as the original model, the objective of a MU method $U$ is to remove the influence of a particular subset of training set $\mathcal{D}_{F}=\{(\mathbf{x}_{i},y_{i})\}_{i=1}^{K}\subset\mathcal{D}$ referred to as the forget set where $K\ll N$. The rest of the training set $\mathcal{D}_{R}=\mathcal{D}\backslash\mathcal{D}_{F}$ is called the retain set. The forget and retain sets are distinct and complementary subsets of the training set. The outcome of MU is an unlearned model $f_{U}$, the aim for which is to perform on par with a model retrained from scratch on $\mathcal{D}_{R}$; this latter model $f_{R}$ is referred to as the retrained model.
We denote the weights of the original model and retrained models’ weights as $\theta_{O}$ and $\theta_{R}$, respectively. For evaluations, we consider two held-out sets: the validation set $\mathcal{D}_{V}$ and test set $\mathcal{D}_{T}$, both drawn from the same distribution as $\mathcal{D}$. We consider the accuracy of the retrained model to be the optimal accuracy. One critical assumption we make is that the MU method has access to $\theta_{O}$, $\mathcal{D}_{R}$, $\mathcal{D}_{F}$, and $\mathcal{D}_{V}$.
3.2 Unlearning Evaluation.
To evaluate the success of unlearning, one approach is to check whether data points in $\mathcal{D}_{F}$ still influence the predictions made by the unlearned model  [44, 14, 45]. This is commonly done via influence functions [17, 46, 47] and membership inference attacks (MIA) [48]. MIA has become one of the most common approaches for evaluating MU methods. It aims to determine whether specific data points were part of the original training dataset based on the unlearned model.
3.2.1 Indiscernibility
@@
We compute the population-based U-MIA, denoted with $\text{MIA}(\mathcal{D},f_{U})$ evaluated on data $\mathcal{D}$ [^1] The methodology for this follows the classic MIA: we compute the losses on $\mathcal{D}_{F}$ and $\mathcal{D}_{V}$, we shuffle and trim them so that they are of equal size. We then train logistic regression models in a 10-fold cross-validation and compute the average accuracy across the folds.. Using this we define the discernibility metric as $\text{Disc}(\mathcal{D}_{V},f_{U})=|2\times\text{MIA}(\mathcal{D},f_{U})-1|\in[0,1]$, and similarly, indiscernibility is given by $\text{Indisc}(\mathcal{D},f_{U})=1-\text{Disc}(\mathcal{D},f_{U})\in[0,1]$. We set $\mathcal{D}$ to either the test set $\mathcal{D}_{T}$ or validation set $\mathcal{D}_{V}$. The indiscernibility equals $1$ when the accuracy of the MIA is not better than random guessing.
A more recent MIA variation is the unlearning likelihood ratio attack (U-LiRA) [14]. As highlighted by [45], U-LiRA is a more robust evaluation approach for approximate MU. Nonetheless, U-LiRA is much more computationally demanding than U-MIA. We evaluate the methods that defeat the weaker, less expensive U-MIA attack, additionally against the U-LiRA attack.
3.2.2 Accuracy
@@
The classification accuracy of the unlearned model on the retain set should be as close as possible to that of the original model. First, we consider three metrics derived directly from the model’s classification accuracy on different sets: retain accuracy (RA), forget accuracy (FA), and test accuracy (TA). RA is defined as
@@
and formulas for FR and TR can be derived using FA and TA, respectively. An unlearned model with a score of $1$ indicates that its accuracy perfectly matches the accuracy of the reference retrained model. A score below $1$ indicates that the model underperforms, and a score above $1$ indicates that the model overperforms.
From these, we can define deviation scores, that provide information on the divergence of the unlearned model from the retrained model’s performance. Namely we consider the retain set deviation $DR$ (Equation 4), forget set deviation $DF$ (Equation 5), and test set deviation $DT$ (Equation 6).
$DR=|RR(f_{U},f_{R})-1|\in[0,+\infty)$ (4)
$DF=|FR(f_{U},f_{R})-1|\in[0,+\infty)$ (5)
$DT=|TR(f_{U},f_{R})-1|\in[0,+\infty)$ (6)
We further define the Retention Deviation (RetDev) as:
$\text{RetDev}=DR+DF+DT\in[0,+\infty),$ (7)
Which provides information on the cumulative divergence of the unlearned model regarding retention score. The closer to 0, the better, as 0 indicates that the model perfectly matches the performance of the retrained model.
3.2.3 Efficiency
run-time efficiency (RTE) of a MU method should ideally be higher than the naive approach of just retraining the model from scratch on the retain set. As the high computational cost of the retraining algorithm motivated the development of approximate MU methods, we evaluate how much faster each MU method is compared to the retraining algorithm.
We define RT of $U$ as the number of seconds it takes to complete, denoted as $\text{RT}(U)$ (we use the same machine and resources for all the experiments). To indicate the relative speedup compared to the retrained model, we define the RTE of an unlearn method $U$ as:
@@ 3.2.3 Efficiency
Where $U_{R}$ denotes the retrain method. The RTE of retraining from scratch is 1; any method with an RTE of less than 1 is slower than retraining from scratch, and vice versa.
@@ 4 Machine Unlearning Methods
Here, we discuss the main unlearning methods considered. First, we cover the classical baselines (Section 4.1). Then, methods from the competition and literature (Section 4.2)
@@ 4.1 Classical Baselines
FineTune (FT) keeps training the original model $f_{O}$ only on the retain set $\mathcal{D}_{R}$ for several epochs; the number of epochs is considered as a hyper-parameter and varies from one dataset and architecture. The models are trained by minimizing the cross-entropy over the retain set.
Successive Random Labels (SRL) trains $f_{O}$ on both the forget set $\mathcal{D}_{F}$ and $\mathcal{D}_{R}$ while the labels of $\mathcal{D}_{F}$ are randomly assigned redrawn from a uniform distribution $\mathcal{U}_{C}$, with $C$ the number of classes at each epoch.
Gradient Ascent (GA) trains the model using gradient ascent steps on the $\mathcal{D}_{F}$. Namely, we compute the cross-entropy loss for a batch and take a step toward the gradient rather than opposite it.
@@ 4.2 State-of-the-art MU methods
In addition to the classical baselines, we evaluate 15 recent MU methods. First, we discuss the seven top-performing methods from the competition. Then, we cover eight methods from the literature. We first discuss the seven top-performing methods from the Machine Unlearning Competition 2023 on Kaggle.
TABLE I: Kaggle - Leaderboard and their associated names in this work.
Kaggle Rank Name Acronym
1 Forget-Contrast-Strengthen (FCS)
2 Masked-Small-Gradients (MSG)
3 Confuse-Finetune-Weaken (CFW)
4 Prune-Reinitialize-Match-Quantize (PRMQ)
5 Convolution-Transpose (CT)
6 Knowledge-Distillation-Entropy (KDE)
7 Repeated-Noise-Injection (RNI)
Forget-Contrast-Strengthen (FCS) [49] minimizes the Kullback-Leibler Divergence (KLD) between the model’s output on $\mathcal{D}_{F}$ and uniform distribution over the output classes, then alternatively optimizes a contrastive loss between the model’s outputs on $\mathcal{D}_{R}$ and $\mathcal{D}_{F}$, and minimizes the cross-entropy loss on $\mathcal{D}_{R}$.